OpenStudy (anonymous):

simplify the expression 6√125a^3b^9c^6 please help !

OpenStudy (annn):

125=25*5

OpenStudy (annn):

a^3=a^2*a

OpenStudy (annn):

b^9=(b^4)^2*b

OpenStudy (annn):

c^6=(c^3)^2

OpenStudy (annn):

do you understand ?

OpenStudy (anonymous):

not really , I don't get any of this